Arizona LBTI Commissioning - 15 October 2010 UT

This wiki page includes John's notes on the center bent Gregorian focal station commissioning aspects of LBTI Commissioning. This represents only a small fraction of the total activity.

Observers: JHill, PHinz, Vidhya, Vanessa, Oli, Paul, Bill, Tom
Telescope Operators: SAllanson, GBechetti
Instrument Support: JMorris
Telescope Support: JLittle, JUrban

Summary

Found a star in each side of LBTI, and overlapped same.

It took a few hours but then there were 22.65 m interference fringes on the star Mira through a 4.5% filter!!!!! At a wavelength of 12.33 microns, these are 124 milliarcsec resolution in imaging mode. See the attached file from Bill Hoffmann below.

The DX Adaptive Secondary still has problems with astigmatism as a function of elevation.

Vidhya's LBTI preset are failing because MCSPU gets an invalid rotator name, or something like that.

The SX acquistion camera for LBTI isn't working correctly.

Preparations

01:20 Update the default LBTI pointing model(s) to the most recent LUCI LFBG pointing model as the starting point.
cp -p LUCIFERSX.20100922.ptmod LBTISX.00000000.ptmod
cp -p LUCIFERSX.20100922.ptmod LBTIDX.00000000.ptmod
We are observing in monocular mode, so the pointing model is the same on both sides.

01:25 Edit PCSInstrument.conf to make LEFTWAVELENGTH and RIGHTWAVELENGTH 2.2 microns to start with. Later LBTI may want a longer wavelength for science observing, or a shorter wavelength for the acquisition camera.

01:30 Copied the present SX LUCI collimation model which matches the above pointing model as the SX LBTI Collimation model. (M3 was already the same.)
cp -p SXPMLUCIFERCollimation.dat SXPMLBTICollimation.dat
cp -p SXSMLUCIFERCollimation.dat SXSMLBTICollimation.dat
01:35 Copied the present DX IRTCF collimation model as the DX LBTI Collimation Model. This collimation model does not match the pointing model.
cp -p DXPMIRTCFCollimation.dat DXPMLBTICollimation.dat
cp -p DXSMIRTCFCollimation.dat DXSMLBTICollimation.dat

Steve sorted out various telescope issues: open-loop error on M3 actuators, problems with DX primary actuator power.

Phil is working on LBTI-MIRAC pupil alignment.

Sky is partly cloudy with loads of cirrus.

Details

Open

04:45 Open to heavy cirrus sky. T=7degC, D=-7degC. Wind 6 m/sec from East.

Then back to zenith for additional instrument work on pupil alignment (on-sky) and MIRAC communications.

05:22 Vidhya sends the first preset. The telescope moves, but the preset fails. We didn't see a star in the 7x7 arcsec field of MIRAC.

05:35 Back to zenith to check some LBTI switches on the acquisition cameras.

06:20 Back to the star......with only the DX acquisition camera working.

Finding the DX star

Found star initiallly on DX acquisition camera at IE=-70 CA=-8 (Acquisition field is 15 arcsec)

Then centered the DX star on MIRAC with IE=-63 CA=-6 (MIRAC field is 7 arcsec in low magnification)

Use eyeball active optics to get image down to a couple arcseconds
Z7=+1000 Z8=-1500 which gives
DX Active X=-1.280 Y=+0.854 Z=-0.8 RX=9.17 RY=13.75
plus a LOAD of -Z6 (-5500 nm in M1 units) (for elevation approximately 60 degrees)

Finding the SX star

Now we run ptspiral to find the SX star on MIRAC (since SX acquisition camera isn't working)
SX IE=-50 CA=+37 Focus +0.21 compared to Zglobal.

Overlapping the images

Now we've detected each of the two images separately. Next we want to bring them together on MIRAC. To match Phil's nominal SX pointing (of LBTI) we want to apply TIp/Tilt to the DX side. SX telescope should also have the better pointing model from LUCI@LFBG.

Tip +13 Tilt +45 didn't work, but approx Tip=-13 Tilt=-45 arcsec applied to DX M2 with the Right PSF Arbitrator brings them close.

Then make a little manual spiral in DX M2 so they overlap at Tip -16 Tilt -41 for EL=58 deg with IE=-50 CA=+37 .

Original Recipe for the overlap

Tip_DX = IE_DX - IE_SX and

Tilt_DX = CA_DX - CA_SX

For tonight we applied the Tip/Tilt to M2 alone (rotating around the coma-neutral point).

It would be better to apply the Tip/Tilt as Mode1 repointing of the telescope (M1M2). The sign of the correction reverses for M1M2 between Tilt and Ry. Chris fixed that in a patch to AO5 and above at the end of October 2010.

Improved Recipe for the overlap (7 November)

If you want to adjust the position by moving DX, use these equations:

Tip_DX = IE_DX + Tip_DX - IE_SX - Tip_SX and

Tilt_DX = CA_DX + Tilt_DX - CA_SX - Tilt_SX

or, if you want to adjust the position by moving SX, use these equations:

Tip_SX = IE_SX + Tip_SX - IE_DX - Tip_DX and

Tilt_SX = CA_SX + Tilt_SX - CA_DX - Tilt_DX

The values on the right side of these equations represent the position where the respective star is found centered on MIRAC after spiraling.

Begin the search for fringes

07:14 Now at 10 microns with Phil overlapping the images, and scanning the phase with LBTI.

When we update the M1 collimation (on both sides) we see the images drift apart in mostly azimuth (up-down on MIRAC). As we might expect, this means the two collimation models have slightly different pointing terms built into them.

08:00 Adjust DX M2 Tip=-16 Tilt=-43, Z=-0.82 Seeing is about 2 arcsec, so eyeball active optics is difficult.

+IE moves the star left on MIRAC display
+CA moves the star down on MIRAC display

Mode 1 M1M2 +Rx moves the star left on MIRAC display
Mode 1 M1M2 +Ry moves the star up on MIRAC display

+Ry is the one that John thinks should be reversed. See Issue 2934.

M2 +Tip moves the star left on the MIRAC display
M2 +Tilt moves the star down on the MIRAC display

M1 +Tip moves the star left on the MIRAC display
M1 +Tilt moves the star down on the MIRAC display

Coordinate Directions (updated April 2012)

Making IE more positive moves the telescope to a lower elevation.
Making CA more positive moves the telescope to a larger azimuth.

On DX CCD47, +AZ is up = +CA
On DX CCD47, +EL is right = -IE

On SX CCD47, +AZ is down = +CA
On SX CCD47, +EL is left = -IE

Recipe for Coma Correction while looking at the MIRAC image

Coma like this: "<" needs +Z7 (positive) sent to PSF
Coma like this: "V" needs +Z8 (positive) sent to PSF

Recipe for Coma Correction while looking at the DX CCD47 image

Coma like this: "<" needs +Z7 (positive) sent to PSF
Coma like this: "V" needs -Z8 (negative) sent to PSF

Changing Targets

08:13 Preset to Mira -- problem with the negative declination not going through correctly on Vidhya's interface.

08:23 Sent the coords with the PRESET GUI. Telescope operators are amazed that the PRESET GUI still works. Adjust IE=-40 CA=+42

Redid the collimation on both sides. Noting that it converges much more quickly when you are looking at the correct star.

09:00 IE=-44

09:13 IE=-46

There are occasional moments of decent seeing where the 10 micron PSFs are nice Airy patterns, but most of the time the PSFs look like birds, dragons, cacti, etc. Seeing appears to be about 2 arcsec.

09:28 The initial scan of pathlength hasn't turned up any fringes. The cirrus has gotten too thick to observe through the 4% filter at 10 microns.

Fringes

09:40 Phil detects a hint of fringes, just after Tom and Paul started down the access road.

09:49 Add -3000 Z6 to DX

09:53 IE=-51 CA=+43

10:00 Fringes have been solidly detected, and recorded on Tom's video camera! Vidhya says "It is so beautiful!".

10:24 John fixes up ptspiral.py so that the command line arguments to set pitch, maxloops and delay work correctly. (This hasn't been applied to the left/right versions yet.)

10:33 Switch to the 9% filter. Clouds are thick enough to ruin any serious data. Vidhya is working on testing guiding.

10:58 Preset to Alpha Gem; On MIRAC at IE=-48 CA=35

Add Z6 +3000? to DX.

Fixed another bug in ptspiral.py where it was reading the wrong value for the starting CA. It was relying on order of the terms in the pointing model.

Preset to Beta Gem Found DX in MIRAC at IE=-52 CA=35 with Tip -16 Tilt -43

Found SX IE=-57 CA=+51 So we moved DX to Tip -16 Tilt -55

Added some +2000 Z7 to SX.

11:50 Vanessa is making WFS measurements on SX with the pyramid.

Sending known amounts of Z7, Z2, Z3, Z8

12:14 Back to Alpha Gem, but the coords were entered incorrectly so we struggled to find it.

Then we had lots of coma. Needed to put +2000 Z7 and -2000 Z8 on M2 as M1 didn't have enough range at EL=80.

Close

12:46 Close for impending dawn and clouds. T=5.6 degC

-- JohnHill - 15 Oct 2010
Topic attachments
I Attachment Action Size Date Who Comment
be909.pngpng be909.png manage 134 K 16 Oct 2010 - 05:19 JohnHill Summary page of first fringes prepared by Bill Hoffmann
pmcc.dx.thermal.20101015.jpgjpg pmcc.dx.thermal.20101015.jpg manage 464 K 18 Oct 2010 - 10:59 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20101015.jpgjpg pmcc.sx.thermal.20101015.jpg manage 489 K 18 Oct 2010 - 11:00 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r15 - 10 Apr 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback